Identify the equation of the parabola with its focus at (-4,7) and the directrix y=1.
A) 24(y-3)=(x+4)^2
B) 12 (y-4)=(x+4)^2
C) 24 (y-7)=(x+4)^2
D) -12(y+4)=(x+4)^2

Answer by Edwin McCravy(20065)

The equation of a parabola is

4p(y-k) = (x-h)^2

Where 

(h,k) is the vertex which is halfway between
the directrix and the focus.

|p| is the distance from the focus to the vertex,
and also from the vertex to the directrix. 

p is taken as a positive number if the parabola opens
upward, i.e., the directrix is below the vertex and focus.

p is taken as a negative number if the parabola opens
downward, i.e., the directrix is above the vertex and focus.  

You can find those and substitute in

4p(y-k) = (x-h)^2

Edwin